# Generalization Error of the Tilted Empirical Risk

### v2

@@ E.2 Information-theoretic Bounds
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma(1-\zeta)\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\end{split}$ (85)
where it holds for $\frac{4I(H;S)\big{(}4\zeta^{-2}-1\big{)}}{\gamma^{2}\kappa_{t}^{2}}\geq n$. Note that, we have,
$\begin{split}&\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}\\ &\geq\frac{|\gamma|}{\zeta}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\\ &\geq\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{split}$ (86)
Therefore, we have,
$\frac{|\gamma|}{\zeta}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\geq\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2},$ (87)
where 87 is a quadratic inequality in $\sqrt{n}$ and using that $(a+b)^{2}\leq 2a^{2}+2b^{2}$ yields,
$\frac{4I(H;S)\big{(}4\zeta^{-2}-1\big{)}}{\gamma^{2}\kappa_{t}^{2}}\geq n.$
As under Assumption E.1 and Lemma C.8, we have $\exp(\gamma\kappa_{t})\leq\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]$, the final result follows. ∎
Using Proposition E.2, we can obtain the following upper bound on the expected generalization error.
Theorem E.4.
Given Assumption E.1, the following upper bound holds on the expected tilted generalization error for $\gamma<0$,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\leq\begin{cases}\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}-\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{-\exp(-\gamma\kappa_{t})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}-\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}\end{cases}\end{split}$
@@
$\begin{split}&\mathbb{E}_{P_{H,S}}[\mathrm{gen}_{\gamma}(H,S)]\\ &=\mathbb{E}_{H,S}[\mathrm{R}(H,\mu)]-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)+\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\\ &\quad+\mathrm{R}_{\gamma}(H,P_{H,S})-\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)].\end{split}$ (88)
Using Lemma C.9, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}[\mathrm{R}(H,\mu)]-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)\leq\frac{|\gamma|}{2}\mathbb{E}_{P_{H}\otimes\mu}[\ell^{2}(H,\tilde{Z})],\end{split}$ (89)
and due to Jensen’s inequality for $\gamma<0$, we have
@@
Using a similar approach as for the proof of Proposition D.9 and applying Proposition E.2, we obtain
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\qquad&\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{\exp(-\gamma\kappa_{t})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\qquad&\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{cases}\end{split}$ (91)
Combining (90), (91) and (89) with (88) completes the proof. ∎
Remark E.5.
Assuming $\gamma=O(n^{-1/2})$, the upper bound in Theorem E.4 has the convergence rate $O(n^{-1/2})$. Note that the result in Theorem E.4 holds for unbounded loss functions, provided that the second moment of the loss function exists.
A similar approach to Theorem E.4 can be applied to derive an information-theoretical lower bound on the expected tilted generalization error for $\gamma<0$ by applying Corollary E.3.
Theorem E.6.
Given Assumption E.1, the following lower bound holds on the expected tilted generalization error for $\gamma<0$,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\geq\begin{cases}-\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}+\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\max\big{(}\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}},\frac{\gamma^{2}\kappa_{t}^{2}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}\big{)}\leq n\\ \frac{\exp(-\gamma\kappa_{t})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}+\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\min(\frac{4\big{(}4\zeta^{-2}-1\big{)}I(H;S)}{\gamma^{2}\kappa_{t}^{2}},\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}})>n.\end{cases}\end{split}$
@@
The proof is similar to Theorem E.4 and using Corollary E.3. ∎
Combining Theorem E.6 and Theorem E.4, we can derive the upper bound on the absolute value of the expected generalization error.
Appendix F Proof and details of Section 5
@@
Under Assumption 4.1 and Assumption 5.1, the difference of the tilted population risk (9) between $\mu$ and $\tilde{\mu}$ is bounded as follows;
$\begin{split}&\Big{|}\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\exp(\gamma\ell(h,\tilde{Z}))])\Big{|}\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})}.\end{split}$ (92)
@@
where (a) and (b) follow from the functional derivative with $\mu_{\lambda}=\tilde{\mu}+\lambda(\mu-\tilde{\mu})$ and Lemma C.16. The same approach can be applied for lower bound. ∎
Proposition 5.3 (Restated).
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$ for $\gamma<0$, then the following upper bound holds on the tilted generalization error $\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&\leq 2\kappa_{s}\exp(-\gamma\kappa_{s})\sqrt{\frac{(\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma}-\frac{\gamma}{2}\kappa_{u}^{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ where $\hat{S}$ is the training dataset under the distributional shift.
@@
where $I_{5}$, $I_{6}$ and $I_{7}$ can be bounded using Lemma C.9, Proposition 5.2 and Theorem 4.2, respectively. ∎
Remark F.1.
It is noteworthy that the upper bound in Proposition 5.3 can be infinite for $\gamma\rightarrow-\infty$ and $\gamma=0$ and a fixed n. Consequently, there must exist a $\gamma\in(-\infty,0)$ that minimizes this upper bound. To illustrate this point, consider the case where $n\rightarrow\infty$; here, the first and second terms in the upper bound would vanish. Thus, we are led to the following minimization problem:
$\gamma^{\star}:=\mathop{\mathrm{arg}\min}_{\gamma\in(-\infty,0)}\left[-\frac{\gamma}{2}\kappa_{u}^{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})}\right],$ (94)
where the solution exists. As $\gamma^{\star}$ decreases when $\mathbb{TV}(\mu,\tilde{\mu})$ increases, practically, this implies that if the training distribution becomes more adversarial (i.e., further away from the benign test distribution), we would use smaller negative $\gamma$’s to bypass outliers.
Proposition 5.4 (Restated).
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, such that the following upper bound holds on the tilted generalization error $\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&\geq-\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}+\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma(1-\zeta)}\\ &\qquad-\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ where $\hat{S}$ is the training dataset under the distributional shift.
@@
where $I_{5}$, $I_{6}$ and $I_{7}$ can be bounded using Lemma C.9, Proposition 5.2 and Theorem 4.3, respectively. ∎
@@
Under the same assumptions in Proposition 5.4, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, and with probability at least $(1-\delta)$, the absolute value of the tilted generalization error under distributional shift satisfies $\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,\hat{S})|&\leq\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{s})B(\delta)}{3n\gamma(1-\zeta)}\\ &\qquad-\frac{\gamma\kappa_{u}^{2}}{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ (95) where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.

### v3

@@ D.2 Proofs and Details of Information-theoretic Bounds under Unbounded Loss
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\leq\frac{-1}{\gamma(1-\zeta^{\prime})\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)},\end{split}$ (63)
Assume for now that there is a $\zeta^{\prime}\in(0,1)$ such that
$\begin{split}&\zeta^{\prime}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}\geq\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}.\end{split}$ (64)
where it holds for $\frac{2I(H;S)}{2\zeta^{\prime}\exp(\gamma\kappa_{t})-|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}\leq n$. Therefore, the final condition is
$\frac{2I(H;S)}{2\zeta^{\prime}\exp(\gamma\kappa_{t})-|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}\leq n\leq\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}$
Note that, we also have due to $2ab\leq a^{2}+b^{2}$,
$\begin{split}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}}\leq\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\end{split}$ (65)
Therefore, we can choose $\zeta=\zeta^{\prime}$. Furthermore, we can also discuss the following cases.
  • If $\zeta\exp(\gamma\kappa_{t})\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}$ holds, then for $n\geq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}$ we have,
    $\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\geq\frac{-\exp(|\gamma|\kappa_{t})}{|\gamma|(1-\zeta)}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}}.\end{split}$
  • If $\zeta\exp(\gamma\kappa_{t})\geq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}$ holds, then we have,
    $\begin{split}&\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\\ &\quad\geq\begin{cases}\frac{-\exp(|\gamma|\kappa_{t})}{|\gamma|(1-\zeta)}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}},\quad&\mbox{ if }\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}\leq n\\ \frac{-\exp(|\gamma|\kappa_{t})}{(1-\zeta)|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)},\quad&\mbox{ if }\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}})>n\geq\frac{2I(H;S)}{2\zeta^{\prime}\exp(\gamma\kappa_{t})-|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}.\end{cases}\end{split}$
Using Proposition 3.9, we can obtain the following upper bound on the expected tilted generalization error.
Proposition D.3.
Given Assumption 3.8, the following upper bound holds on the expected tilted generalization error for $\gamma<0$,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\leq\begin{cases}\frac{\exp(|\gamma|\kappa_{t})}{|\gamma|}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}}+|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon},\qquad&\mbox{if }\frac{I(H;S)}{n}\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\\ \frac{\exp(|\gamma|\kappa_{t})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)}+|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon},\qquad&\mbox{ if }\frac{I(H;S)}{n}>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\end{cases}\end{split}$
@@
$\mathbb{E}_{P_{H,S}}[\mathrm{gen}_{\gamma}(H,S)]$ $=\mathbb{E}_{P_{H,S}}[\mathrm{R}(H,\mu)]-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)+\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})$ (66)
$\quad+\mathrm{R}_{\gamma}(H,P_{H,S})-\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)].$
Then, using Lemma C.9, we have
$\mathbb{E}_{P_{H,S}}[\mathrm{R}(H,\mu)]-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)\leq|\gamma|^{\epsilon}\mathbb{E}_{P_{H}\otimes\mu}[\ell^{1+\epsilon}(H,\tilde{Z})]\leq|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon},$ (67)
and now Jensen’s inequality for $\gamma<0$ yields
@@
Applying Proposition 3.9, we obtain
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\frac{\exp(|\gamma|\kappa_{t})}{|\gamma|}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\\ \frac{\exp(|\gamma|\kappa_{t})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)},\qquad&\mbox{if }\frac{I(H;S)}{n}>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\end{cases}\end{split}$
Combining (68), (D.2) and (67) with (66) completes the proof. ∎
Proposition D.4.
Under Assumption 3.8, there exists $\zeta\in(0,1)$ such that one of the following cases holds,
(a) $\zeta\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $n\geq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}$,
(b) $\zeta>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $n\geq\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}$,
(c) $\zeta>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}>n\geq\frac{2I(H;S)}{2\zeta\exp(\gamma\kappa_{t})-|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}$,
Then, the following lower bounds on expected tilted generalization error hold,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)\geq\begin{cases}\frac{-\exp(|\gamma|\kappa_{t})}{|\gamma|(1-\zeta)}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}}-|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon},\mbox{ \text{if (a) or (b) hold,}}\\ \frac{-\exp(|\gamma|\kappa_{t})}{(1-\zeta)|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)}-|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon},\mbox{ \text{if (c) holds}}.\end{cases}\end{split}$
@@
The proof is similar to Proposition D.3 and using Proposition 3.10. ∎
Theorem 3.11.
Under Assumption 3.8, there exists $\zeta\in(0,1)$ such that one of the same cases in (a) $\zeta\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $n\geq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}$, (b) $\zeta>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $n\geq\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}$, (c) $\zeta>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}>n\geq\frac{2I(H;S)}{2\zeta\exp(\gamma\kappa_{t})-|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}$, Then, the following lower bounds on non-linear expected generalization error hold, $\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|\leq\begin{cases}&\frac{\exp(|\gamma|\kappa_{t})}{|\gamma|(1-\zeta)}\sqrt{\frac{2\kappa_{t}^{2}I(H;S)}{n}}+|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon},\mbox{ \text{if (a) or (b) hold,}}\\ &\frac{\exp(|\gamma|\kappa_{t})}{(1-\zeta)|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)}+|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon},\mbox{ \text{if (c) holds}}.\end{cases}\end{split}$
Proof.
This result follows by combining the upper bound from Proposition D.3 with the lower bound from Proposition D.4. ∎
Remark D.5 (Individual Sample Bound Discussion).
We can derive the results based on individual sample approach (Bu et al., 2020), replacing the following inequality with (53), which is based on individual sample,
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}\sqrt{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\frac{1}{n}\sum_{i=1}^{n}I(H;Z_{i})}&\quad\mbox{ if }\max_{i\in[n]}I(H;Z_{i})\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\\ \frac{1}{n}\sum_{i=1}^{n}I(H;Z_{i})+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}&\mbox{ if }\min_{i\in[n]}I(H;Z_{i})>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}.\end{cases}\end{split}$ (69)
Appendix E Proof and details of Section 4
@@
Given Assumption 4.1 and Assumption 3.1, then the difference of tilted population risk under, (9), between $\mu$ and $\tilde{\mu}$ is bounded as follows for all $h\in\mathcal{H}$,
$\begin{split}&\Big{|}\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\exp(\gamma\ell(h,\tilde{Z}))])\Big{|}\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\Big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\Big{)}}{(\kappa_{u}-\kappa_{s})}.\end{split}$ (70)
@@
where (a) and (b) follow from the functional derivative with $\mu_{\lambda}=\tilde{\mu}+\lambda(\mu-\tilde{\mu})$ and Lemma C.16. The same approach can be applied for the lower bound. ∎
Remark E.1.
For positive $\gamma$, the result in Proposition 4.2 does not hold and can be unbounded.
Proposition E.2 (Upper Bound).
Given Assumptions  3.1 and 4.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$ for $\gamma<0$, then the following upper bound holds on the tilted generalization error $\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&\leq\frac{2\exp(|\gamma|\kappa_{s})}{|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{s}^{1+\epsilon}\log(2/\delta)}{n}}\\ &\qquad+\frac{4\exp(|\gamma|\kappa_{s})(\log(2/\delta))}{3n|\gamma|}+|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ where $\hat{S}$ is the training dataset under the distributional shift.
@@
where $I_{5}$, $I_{6}$ and $I_{7}$ can be bounded using Lemma C.9, Proposition 4.2 and Proposition 3.3, respectively. ∎
Proposition E.3 (Lower Bound).
Given Assumptions  3.1 and 4.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, such that the following upper bound holds on the tilted generalization error $\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})\geq&-\frac{2\exp(|\gamma|\kappa_{s})}{(1-\zeta)|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{s}^{1+\epsilon}\log(2/\delta)}{n}}-\frac{4\exp(|\gamma|\kappa_{s})(\log(2/\delta))}{3n|\gamma|(1-\zeta)}\\ &\quad\quad-\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ where $\hat{S}$ is the training dataset under the distributional shift.
@@
where $I_{5}$, $I_{6}$ and $I_{7}$ can be bounded using Lemma C.9, Proposition 4.2 and Proposition 3.4, respectively. ∎
@@
Under the same assumptions in Theorem 3.5, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, and with probability at least $(1-\delta)$, the absolute value of the tilted generalization error under distributional shift satisfies $\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,\hat{S})|$ (72) $\leq\frac{2\exp(|\gamma|\kappa_{u})}{(1-\zeta)|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}B(\delta)}{n}}+\frac{4\exp(|\gamma|\kappa_{u})B(\delta)}{3n|\gamma|(1-\zeta)}$ $+|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},$ where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.